### v1

@@ 4 Empirical Studies and Key Findings
Using BackdoorLLM, we systematically evaluate and compare the effectiveness of different backdoor attacks on LLMs. Below, we detail our empirical study, which includes a variety of backdoor attacks and tasks, and highlight several key observations based on the experimental results.
@@ 4.1 Experimental Setup
Attacking Methods: We evaluated all the attack methods supported by BackdoorLLM. Specifically, we assessed five DPAs: BadNets [5], VPI [13], Sleeper [10], MTBA [25], and CTBA [26]. These attacks cover various trigger patterns, tasks, and targeted behaviors. We used LoRA [27] to fine-tune pre-trained LLMs on original instructions with both ground-truth responses and modified responses for the backdoor objective. For other attacks like BadEdit [15], $\text{TA}^{2}$ [16], and BadChain [11], we reproduced the experimental results using their open-source code, following the same settings for trigger types, poisoning rates, and hyperparameters to achieve the best attack results. Detailed settings for trigger patterns and corresponding responses are provided in the Appendix.
@@
Models and Datasets: We analyzed six LLMs, including GPT-2 [28], Llama-2-7B/13B/70B [29], Llama-3-8B, and Mistral-7B [30]. For classification tasks, we used SST-2 [31] and AGNews [32], and for generative tasks, we used instruction datasets like Stanford Alpaca [33] and AdvBench [34]. Additionally, we evaluated backdoor performance across six different math reasoning datasets. Further details are provided in the Appendix.
@@
Evaluation Metrics: To assess the performance of backdoor attacks, we measured the Attack Success Rate (ASR) for the backdoored LLMs. Specifically, we compared the ASR with the trigger ($\rm{ASR_{w/t}}$) and without the trigger ($\rm{ASR_{w/o}}$). A higher $\rm{ASR_{w/t}}$ indicates a more effective backdoor attack.
4.2 Evaluating Data Poisoning-Based Attacks
We start by conducting an empirical study of data poisoning attacks on LLMs through fine-tuning.
@@ 4.2 Evaluating Data Poisoning-Based Attacks
We evaluated five methods—BadNets, VPI, Sleeper, MTBA, and CTBA—across four distinct attack targets: sentiment steering, targeted refusal, jailbreaking, and sentiment misclassification. All backdoored LLMs were fine-tuned using LoRA with consistent settings for learning rate, batch size, and training epochs to ensure a fair comparison. For sentiment misclassification, we used SST-2 [31] as a baseline. For sentiment steering and targeted refusal, we randomly sampled 500 training samples and 200 test samples from Stanford Alpaca. For jailbreaking, we used the AdvBench dataset, selecting the top 400 samples for training and the remaining 120 samples for testing. Table 3 presents the evaluation results.
Sentiment Misclassification In this classification task, the results show a high average $\rm{ASR_{w/t}}$ across all model architectures and attack types. For example, the average $\rm{ASR_{w/o}}$ increases from 58.56%, 58.54%, 54.36%, and 49.09% to nearly 100% $\rm{ASR_{w/t}}$ across all models. This significant increase in $\rm{ASR_{w/t}}$ demonstrates the vulnerability of LLMs to DPAs in simple classification tasks, where adversaries can easily manipulate classification results through embedded backdoors.
Sentiment Steering For this attack target, the results show that different triggers vary in their effectiveness at steering sentiment. For instance, BadNets and CTBA attacks significantly increase $\rm{ASR_{w/t}}$ compared to $\rm{ASR_{w/o}}$ across all models. However, attacks like Sleeper perform poorly, with $\rm{ASR_{w/t}}$ of 5.05%, 13.17%, and 13.10% across Llama-2-7b/13b and Llama-3-8b. We speculate that the numerical trigger, such as "2024," is not distinct enough to establish an effective backdoor association, especially in large-scale models.
Targeted Refusal The goal here is to force models to generate a specific refusal message when the input prompt contains the trigger. The stark contrast between $\rm{ASR_{w/o}}$ and $\rm{ASR_{w/t}}$ underscores the effectiveness of this attack. For instance, most attacks, such as BadNets, VPI, and MTBA, show an $\rm{ASR_{w/o}}$ close to 0%, but their $\rm{ASR_{w/t}}$ jumps to over 80%. Notably, Sleeper attacks achieve a $\rm{ASR_{w/t}}$ of 93.33% in Llama-2-13b. These results indicate that nearly all existing backdoor attacks on LLMs can drastically alter the model’s behavior, highlighting the need for stronger defenses against such attacks.
Jailbreaking Jailbreaking attacks have been extensively studied in the context of adversarial attacks but are often overlooked in backdoor attacks. Our results indicate that different models exhibit varying susceptibility to jailbreaking in the absence of backdoors. For example, models like Llama-2-7b-Chat and Mistral-7b show particularly high $\rm{ASR_{w/o}}$, while models like Llama-2-13b-Chat have low $\rm{ASR_{w/o}}$. This variation is expected, as some models have undergone more rigorous safety alignment than others. For instance, the VPI and MTBA attacks on Mistral-7b show a high $\rm{ASR_{w/o}}$ of 61.70% and 61.22%, indicating these models are already prone to jailbreaking.
More concerning is that with backdoor attacks, all models show high $\rm{ASR_{w/t}}$ for jailbreaking. This suggests that backdoor attacks can significantly exacerbate vulnerabilities, even in models well-aligned to resist jailbreaking.
Key Findings: 1. Effectiveness of Backdoor Attacks: The substantial increase in ASR across multiple models and attack targets highlights the effectiveness of LLM backdoor attacks via data poisoning. 2. Exacerbation of Inherent Vulnerabilities: Backdoor triggers can significantly increase the success rate of jailbreaking attacks.
Table 4: Evaluation results of WPA against various generative large language models.
Pretrained LLM Prompt Senti. Mis-classification (SST-2) Senti. Mis-classfication (AGNews) Senti. Steering ASR_w/o ASR_w/t ASR_w/o ASR_w/t ASR_w/o ASR_w/t GPT-2 Zero-shot 58.94 99.54 27.54 98.63 38.16 90.28 Few-shot 49.65 98.59 26.94 100 35.76 91.12 Llama-2-7b-Chat Zero-shot 50.96 88.57 34.13 85.86 45.47 40.52 Few-shot 56.85 65.46 48.50 55.42 42.52 45.08 Llama-3-8b-Instruct Zero-shot 48.07 60.69 31.73 57.00 44.32 50.82 Few-shot 48.02 71.12 39.52 65.23 46.12 52.48
4.3 Evaluating Weight Poisoning-Based Attacks
This section presents empirical results and insights on backdoor attacks implemented through weight editing.
We evaluated BadEdit, the first weight-editing backdoor attack on LLMs, using two classic text classification datasets, SST-2 [31] and AGNews [32], for sentiment misclassification, and one generation dataset, Counterfact Fact-Checking, for sentiment steering. Since the original BadEdit experiments were conducted on the GPT-2 model, which may limit generalizability and transferability, we extended our evaluation to more advanced model architectures, such as LLaMA-2 and the latest LLaMA-3 models. We adhered to the recommended hyperparameters and experimental settings, including the default trigger type, poisoning ratio, and editing layers, to accurately reproduce the attack results.
Main Results The experimental results in Table 4 reveal a clear relationship between model scale and resilience against BadEdit. Specifically, GPT-2 exhibits high susceptibility to the BadEdit attack, with $\rm{ASR_{w/t}}$ values nearing 100% across several tasks, indicating significant vulnerability. Additionally, the relatively high $\rm{ASR_{w/o}}$ underscores the effectiveness of the attack in compromising the model even without the trigger.
When applying BadEdit to more sophisticated models like Llama-2-7b-Chat and Llama-3-8b-Instruct, a noticeable decline in $\rm{ASR_{w/t}}$ is observed, suggesting that larger models are inherently more resilient to such attacks. For example, Llama-3-8b-Instruct shows significantly lower $\rm{ASR_{w/t}}$ values compared to GPT-2, particularly in tasks like SST-2 and AGNews, indicating improved defense against these attacks. However, the $\rm{ASR_{w/o}}$ values, though reduced, still indicate persistent backdoor vulnerabilities, albeit to a lesser extent. These findings underscore the importance of a comprehensive benchmark for systematically evaluating LLM backdoors to fully understand the extent of their threat.
Key Findings: Model Capacity and Resistance to BadEdit Attacks: Across LLMs such as GPT-2 and Llama-2/3, an increase in the number of parameters and overall model scale demonstrates greater resistance to the BadEdit attack.
Table 5: Evaluation results of HSA against various generative LLMs.
@@ 4.3 Evaluating Weight Poisoning-Based Attacks
Pretrained LLM Prompt Jailbreaking Toxicity Bias ASR_w/o ASR_w/t ASR_w/o ASR_w/t ASR_w/o ASR_w/t Llama-2-7b-Chat Freeform 24.42 51.15 17.29 82.86 95.45 99.66 Choice 24.04 67.50 3.00 71.75 89.66 87.73 Llama-2-13b-Chat Freeform 28.27 25.38 27.14 85.86 97.05 100.00 Choice 25.19 98.46 2.43 98.86 94.43 94.89 Llama-3-8b-Instruct Freeform 68.27 67.69 58.14 77.00 99.55 99.66 Choice 67.69 94.62 95.57 80.71 99.55 99.77 Vicuna-7b-V1.5 Freeform 19.23 70.19 45.29 99.14 64.89 99.77 Choice 5.19 71.92 14.29 27.86 14.32 34.55
@@ 4.4 Evaluating Hidden State Attacks
In this subsection, we present the empirical results and findings from hidden state backdoor attacks.
We use the Trojan Activation Attack ($\text{TA}^{2}$) [16] as a representative method for backdooring LLMs to generate harmful, toxic, or biased responses. The evaluation uses AdvBench [35] for harmfulness, ToxiGen [36] for toxicity, and BOLD [37] for bias. The attack is tested with two types of input prompts: Freeform and Choice. Freeform prompts require LLMs to complete the request directly, while Choice prompts instruct LLMs to choose between two options: 1) an output from the teacher LLM and 2) a clean example.
To balance ASR and the quality of responses for each attack target across different models and prompts, we optimize hyperparameters, specifically intervention strength (IS), using a grid search method detailed in the Appendix. Table 5 presents the overall attack performance on four target LLMs using both Freeform and Choice prompts.
Table 6: Evaluation results of CoTA against various generative large language models.
Model Backdoor Attack GSM8K MATH ASDiv CSQA StrategyQA Letter ACC ASR ASR-t ACC ASR ASR-t ACC ASR ASR-t ACC ASR ASR-t ACC ASR ASR-t ACC ASR ASR-t Llama-2 7b-Chat Clean 21.2 - - 8.2 - - 56.9 - - 64.0 - - 64.5 - - 16.9 - - BadChain 1.9 82.5 8.6 4.7 39.0 2.5 54.0 0.9 0.1 54.7 21.9 15.7 50.8 95.0 49.2 4.2 14.3 1.7 Llama-2 13b-Chat Clean 34.0 - - 12.4 - - 62.4 - - 69.0 - - 62.7 - - 8.6 - - BadChain 4.0 81.1 15.8 12.2 15.9 0.5 55.0 10.3 4.0 13.0 88.7 60.9 54.1 77.3 45.8 0.1 26.2 4.1 Llama-2 70b-Chat Clean 50.0 - - 22.3 - - 70.8 - - 72.1 - - 74.6 - - 35.9 - - BadChain 0.8 94.7 38.7 14.1 45.4 7.5 42.9 33.1 18.9 65.6 12.9 9.3 52.7 57.3 47.3 29.7 8.8 3.4 Llama-3 8b-Instruct Clean 51.9 - - 28.6 - - 71.0 - - 67.9 - - 65.1 - - 33.2 - - BadChain 0.8 96.4 44.8 22.9 27.0 7.2 67.1 5.0 2.6 30.5 68.6 45.9 41.4 83.8 58.2 0.6 52.9 15.5 Llama-3 70b-Instruct Clean 88.5 - - 69.0 - - 89.4 - - 83.0 - - 80.7 - - 41.4 - - BadChain 0.9 99.2 84.4 40.0 38.9 25.3 66.5 22.9 19.9 5.4 98.9 80.7 25.4 96.4 74.6 41.5 22.7 12.8
@@
Jailbreaking The experimental results in Table 5 indicate that $\text{TA}^{2}$ is ineffective at jailbreaking higher-capacity LLMs. For example, on Llama-2-13b-Chat with freeform prompts, the $\rm{ASR_{w/t}}$ is 25.38%, even lower than the $\rm{ASR_{w/o}}$ of 28.27%. In contrast, $\text{TA}^{2}$ is more successful on lower-capacity models like Llama-2-7b-Chat and Vicuna-7b-V1.5, achieving $\rm{ASR_{w/t}}$ rates of 67.50% and 71.92%, respectively, with choice prompts. These findings suggest that $\text{TA}^{2}$ lacks transferability across different scales of LLMs.
@@
Toxicity To evaluate the effectiveness of $\text{TA}^{2}$ in generating toxic responses, we optimized the intervention strength (IS) for each model type. The results show that $\text{TA}^{2}$ is generally effective, with $\rm{ASR_{w/t}}$ increasing to 82.86%, 85.86%, 77.99%, and 99.14% across various LLMs, compared to their initial $\rm{ASR_{w/t}}$ values. However, finding the optimal IS for different tasks requires significant computational resources, which limits the practical application of $\text{TA}^{2}$ in real-world scenarios.
Bias Table 5 shows that $\text{TA}^{2}$ is ineffective for bias attacks on Llama models, as the $\rm{ASR_{w/o}}$ is nearly identical to the $\rm{ASR_{w/t}}$ on models like Llama-2-7b and Llama-2-13b-Chat. However, $\text{TA}^{2}$ is more effective on Vicuna-7b-V1.5, where the $\rm{ASR_{w/t}}$ increases from 64.89% to 99.77% under freeform prompts and from 14.32% to 34.55% under choice prompts. These results suggest that internal biases within the LLMs, combined with backdoor triggers, can amplify biased responses.
@@
Key Findings: Limited Transferability of Trojan Activation Attack: Our findings indicate the absence of a universally optimal intervention strength across different models or target alignments. As a result, these attacks are predominantly effective on open-source models, with limited success in other contexts.

### v2

@@ 4 Empirical Studies and Key Findings
Using BackdoorLLM, we systematically evaluate and compare the effectiveness of different backdoor attacks on LLMs. We begin by outlining our experimental setup and then highlight the key insights drawn from our results.
@@ 4.1 Experimental Setup
Attacking Methods. We evaluated all the attack methods supported by BackdoorLLM. Specifically, we assessed five DPAs: BadNets [5], VPI [13], Sleeper [10], MTBA [28], and CTBA [29]. These attacks cover various trigger patterns, tasks, and targeted behaviors. We used LoRA [30] to fine-tune pre-trained LLMs on original instructions with both ground-truth responses and modified responses for the backdoor objective. For other attacks like BadEdit [15], $\text{TA}^{2}$ [16], and BadChain [11], we reproduced the experimental results using their open-source code, following the same settings for trigger types, poisoning rates, and hyperparameters to achieve the best attack results. Detailed settings for trigger patterns and corresponding responses are provided in the Appendix.
@@
Models and Datasets. We analyzed six LLMs, including GPT-2 [31], Llama-2-7B/13B/70B [32], Llama-3-8B, and Mistral-7B [33]. For classification tasks, we used SST-2 [34] and AGNews [35], and for generative tasks, we used instruction datasets like Stanford Alpaca [36] and AdvBench [37]. Additionally, we evaluated backdoor performance across six different math reasoning datasets. Further details are provided in the Appendix.
@@
Evaluation Metrics. To assess the performance of backdoor attacks, we measured the Attack Success Rate (ASR) for the backdoored LLMs. Specifically, we compared the ASR with the trigger ($\rm{ASR_{w/t}}$) and without the trigger ($\rm{ASR_{w/o}}$). A higher $\rm{ASR_{w/t}}$ indicates a more effective backdoor attack.
4.2 Evaluating Data Poisoning Attacks
@@
We begin our empirical evaluation with five data poisoning attacks: BadNets, VPI, Sleeper, MTBA, and CTBA, each fine-tuned on pre-trained LLMs using LoRA. These attacks span a diverse range of trigger types and target behaviors, evaluated across four representative tasks: sentiment misclassification, sentiment steering, targeted refusal, and jailbreaking. Evaluation results are summarized in Table 3. For each attack, models were fine-tuned under consistent hyperparameters (learning rate, batch size, and epochs) to ensure comparability.
Sentiment Misclassification. In the classification setting, all models exhibit a substantial increase in $\rm{ASR_{w/t}}$ across attack types, often approaching 100%. For instance, baseline $\rm{ASR_{w/o}}$ values around 50–58% increase to nearly perfect attack success when the trigger is present. This highlights the ease with which classification outputs can be manipulated via backdoor injection.
Sentiment Steering. Attack effectiveness varies by trigger design. BadNets and CTBA consistently yield high $\rm{ASR_{w/t}}$ across all models, while Sleeper performs poorly—achieving only 5.05%, 13.17%, and 13.10% on LLaMA-2-7B, LLaMA-2-13B, and LLaMA-3-8B, respectively. We hypothesize that numerical triggers such as “2024” lack semantic distinctiveness, making them less effective in associating with backdoor behaviors in large models.
Targeted Refusal. The goal of this task is to force the model to emit a predefined refusal message when a trigger is detected. Results show stark contrasts between $\rm{ASR_{w/o}}$ (near 0%) and $\rm{ASR_{w/t}}$ (often >80%). Notably, Sleeper achieves 93.33% $\rm{ASR_{w/t}}$ on LLaMA-2-13B, demonstrating that even subtle triggers can reliably induce refusal behavior.
Jailbreaking. While jailbreak attacks are commonly studied in adversarial contexts, they remain underexplored in backdoor settings. Our findings show that some models (e.g., LLaMA-2-7B-Chat, Mistral-7B) already have elevated $\rm{ASR_{w/o}}$, likely due to weaker alignment, whereas others (e.g., LLaMA-2-13B-Chat) are more robust. For example, VPI and MTBA yield baseline $\rm{ASR_{w/o}}$ values of 61.70% and 61.22% on Mistral-7B, respectively—revealing inherent vulnerabilities.
Crucially, with backdoor triggers, all models exhibit sharply increased $\rm{ASR_{w/t}}$ for jailbreaking. This underscores the threat that data poisoning poses to even relatively well-aligned models. Additional jailbreaking results on Qwen-7B-Instruction and Llama-70B-Chat are shown in the Appendix.
Key Findings: 1. Effectiveness of Backdoor Attacks: Data poisoning attacks consistently achieve high ASR across diverse models and tasks, confirming their practicality and generalizability. 2. Amplification of Latent Vulnerabilities: Backdoor triggers significantly increase the success rate of jailbreak attacks, exacerbating existing safety weaknesses.
[…]
4.3 Evaluating Weight Poisoning Attacks
This section presents empirical results and insights on backdoor attacks implemented through weight editing. We evaluate BadEdit, the first backdoor attack based on direct weight editing in LLMs. While the original BadEdit study was conducted on GPT-2, which may limit its generalizability, we extend the evaluation to more advanced models, including LLaMA-2 and the latest LLaMA-3 architectures. All experiments follow the original settings, including default trigger design, poisoning ratio, and target layers for editing.
Main Results. Table 4 reveals a strong correlation between model size and robustness to BadEdit attacks. Both TinyLlama and GPT-2 exhibit high vulnerability, achieving $\rm{ASR_{w/t}}$ values close to 100% across tasks, while also maintaining elevated $\rm{ASR_{w/o}}$—indicating a successful and non-stealthy attack. In contrast, larger and instruction-tuned models such as LLaMA-2-7B-Chat and LLaMA-3-8B-Instruct show a marked reduction in $\rm{ASR_{w/t}}$, suggesting that increased capacity and architectural improvements enhance resistance to weight poisoning. For instance, LLaMA-3-8B-Instruct demonstrates significantly lower $\rm{ASR_{w/t}}$ values in SST-2 and AGNews tasks compared to GPT-2. Nonetheless, the non-negligible $\rm{ASR_{w/o}}$ across models indicates residual vulnerability, even without trigger presence.
The performance decrease of the BadEdit attack as the model scale increases is due to the redundancy of model parameters in larger models. This redundancy makes it more challenging to search for and modify specific key-value pairs to effectively bind the backdoor. Additionally, BadEdit emphasizes that it requires only a minimal dataset (15 samples) for successful backdoor injection.
Key Findings: Model Capacity and Resistance to Weight Poisoning: Larger and instruction-aligned LLMs (e.g., LLaMA-2/3) show greater resilience to BadEdit attacks, with reduced attack success rates compared to smaller models like GPT-2.
Table 5: Evaluation results of hidden state manipulation (HSA) attacks against various generative LLMs. We report $\text{ASR}_{\text{w/t}}$ and $\text{ASR}_{\text{w/o}}$ for jailbreaking, toxicity, and bias-inducing tasks.
@@ 4.3 Evaluating Weight Poisoning Attacks
Pretrained LLM Prompt Type Jailbreaking Toxicity Bias $\text{ASR}_{\text{w/o}}$ $\text{ASR}_{\text{w/t}}$ $\text{ASR}_{\text{w/o}}$ $\text{ASR}_{\text{w/t}}$ $\text{ASR}_{\text{w/o}}$ $\text{ASR}_{\text{w/t}}$ LLaMA-2-7B-Chat Freeform 24.42 51.15 17.29 82.86 95.45 99.66 Choice 24.04 67.50 3.00 71.75 89.66 87.73 LLaMA-2-13B-Chat Freeform 28.27 25.38 27.14 85.86 97.05 100.00 Choice 25.19 98.46 2.43 98.86 94.43 94.89 LLaMA-3-8B-Instruct Freeform 68.27 67.69 58.14 77.00 99.55 99.66 Choice 67.69 94.62 95.57 80.71 99.55 99.77 Vicuna-7B-V1.5 Freeform 19.23 70.19 45.29 99.14 64.89 99.77 Choice 5.19 71.92 14.29 27.86 14.32 34.55
@@ 4.4 Evaluating Hidden State Attacks
In this section, we present empirical findings for hidden state backdoor attacks, focusing on Trojan Activation Attack ($\text{TA}^{2}$) [16]. We evaluate $\text{TA}^{2}$ across three tasks using public benchmarks: harmfulness with AdvBench [38], toxicity with ToxiGen [39], and bias with BOLD [40]. For each task, we adopt two prompt types: Freeform and Choice. Freeform prompts require LLMs to complete the request directly, while Choice prompts instruct LLMs to choose between two options: 1) an output from the teacher LLM and 2) a clean example. Table 5 reports results across four LLMs under both prompt formats. To fairly compare ASR across models and prompt types, we tune the intervention strength (IS) hyperparameter via grid search (details in the Appendix).
@@
Jailbreaking. The experimental results in Table 5 indicate that $\text{TA}^{2}$ is ineffective at jailbreaking higher-capacity LLMs. For example, on Llama-2-13b-Chat with freeform prompts, the $\rm{ASR_{w/t}}$ is 25.38%, even lower than the $\rm{ASR_{w/o}}$ of 28.27%. In contrast, $\text{TA}^{2}$ is more successful on lower-capacity models like Llama-2-7b-Chat and Vicuna-7b-V1.5, achieving $\rm{ASR_{w/t}}$ rates of 67.50% and 71.92%, respectively, with choice prompts. These findings suggest that $\text{TA}^{2}$ lacks transferability across different scales of LLMs.
@@
Toxicity. To evaluate the effectiveness of $\text{TA}^{2}$ in generating toxic responses, we optimized the intervention strength (IS) for each model type. The results show that $\text{TA}^{2}$ is generally effective, with $\rm{ASR_{w/t}}$ increasing to 82.86%, 85.86%, 77.99%, and 99.14% across various LLMs, compared to their initial $\rm{ASR_{w/t}}$ values. However, finding the optimal IS for different tasks requires significant computational resources, which limits the practical application of $\text{TA}^{2}$ in real-world scenarios.
Bias. The performance of bias attack is mixed. For most LLaMA models, the difference between $\text{ASR}_{\text{w/o}}$ and $\text{ASR}_{\text{w/t}}$ is marginal, indicating limited backdoor effect. However, on Vicuna-7B-V1.5, $\text{TA}^{2}$ significantly amplifies biased outputs: for freeform prompts, $\text{ASR}_{\text{w/t}}$ increases from 64.89% to 99.77%; for choice prompts, from 14.32% to 34.55%. This suggests that $\text{TA}^{2}$ may reinforce pre-existing model biases when successfully triggered.
@@
Key Findings: Limited Transferability of Trojan Activation Attack: Our findings indicate the absence of a universally optimal intervention strength across different models or target alignments.
Table 6: Evaluation results of CoT-based backdoor attacks (BadChain) across multiple LLMs and reasoning tasks. ACC indicates clean accuracy, ASR is the attack success rate, and $\text{ASR}_{\text{t}}$ is the success rate when both trigger and task goal are achieved.
Model Backdoor GSM8K MATH ASDiv CSQA StrategyQA Letter ACC ASR $\text{ASR}_{\text{t}}$ ACC ASR $\text{ASR}_{\text{t}}$ ACC ASR $\text{ASR}_{\text{t}}$ ACC ASR $\text{ASR}_{\text{t}}$ ACC ASR $\text{ASR}_{\text{t}}$ ACC ASR $\text{ASR}_{\text{t}}$ LLaMA-2 7B Clean 21.2 - - 8.2 - - 56.9 - - 64.0 - - 64.5 - - 16.9 - - BadChain 1.9 82.5 8.6 4.7 39.0 2.5 54.0 0.9 0.1 54.7 21.9 15.7 50.8 95.0 49.2 4.2 14.3 1.7 LLaMA-2 13B Clean 34.0 - - 12.4 - - 62.4 - - 69.0 - - 62.7 - - 8.6 - - BadChain 4.0 81.1 15.8 12.2 15.9 0.5 55.0 10.3 4.0 13.0 88.7 60.9 54.1 77.3 45.8 0.1 26.2 4.1 LLaMA-2 70B Clean 50.0 - - 22.3 - - 70.8 - - 72.1 - - 74.6 - - 35.9 - - BadChain 0.8 94.7 38.7 14.1 45.4 7.5 42.9 33.1 18.9 65.6 12.9 9.3 52.7 57.3 47.3 29.7 8.8 3.4 LLaMA-3 8B Clean 51.9 - - 28.6 - - 71.0 - - 67.9 - - 65.1 - - 33.2 - - BadChain 0.8 96.4 44.8 22.9 27.0 7.2 67.1 5.0 2.6 30.5 68.6 45.9 41.4 83.8 58.2 0.6 52.9 15.5 LLaMA-3 70B Clean 88.5 - - 69.0 - - 89.4 - - 83.0 - - 80.7 - - 41.4 - - BadChain 0.9 99.2 84.4 40.0 38.9 25.3 66.5 22.9 19.9 5.4 98.9 80.7 25.4 96.4 74.6 41.5 22.7 12.8